# Inferring Quantum Network Topologies using Genetic Optimisation of Indirect Measurements

### v2

@@ I Introduction
Quantum networks provide a powerful platform for distributing, processing, and storing quantum information across multiple interconnected systems, playing a significant role in the advancement of emerging quantum technologies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Reconstructing the topology of unknown networks is therefore a fundamental task for quantum technological applications. Continuous-time quantum walks (CTQWs) [14, 15] are an effective tool for modelling the topology-driven dynamics in complex quantum networks. They have been successfully used to investigate the dynamics of transport phenomena in numerous systems [16, 17]. A key example is the demonstration of efficient energy transfer in the Fenna-Matthews-Olson complex [18], one of the most extensively studied light-harvesting protein-pigment complexes [19, 20, 21]. The complete characterisation of a network’s topology can be achieved using methods such as quantum process tomography [22], which requires however direct measurement of the state of the entire network, with the number of required measurements scaling exponentially with the system’s dimension. As a result, such an approach quickly becomes infeasible for even moderate network sizes [23]. Approaches based on machine learning techniques [24, 25, 26] provide alternative methods for topology reconstruction with improved measurement scaling, they are however limited to case-specific scenarios and have not been trialled beyond moderate system sizes.
@@
This work tackles a fundamental challenge of inferring the topology of an unknown quantum network [27] by leveraging indirect system measurements using a probe. This approach provides a non-invasive framework for interacting with complex quantum systems without the need for direct access to the network itself. Such probing procedure consists of injecting an excitation into a network node and attaching an external sink to another one, which we model as an irreversible decay channel. The excitation crosses the network to decay into the sink. By measuring the sink, we collect data driven by the underlying network topology. We thus seek to exploit this relationship to infer the topology of the network.
@@
We frame the task as an optimization problem with the goal of constructing a network with measurement data as similar to the target network as possible. As the parameters of this problem are the connections between network nodes, the parameter space is discretised and so lends itself to an approach such as a genetic algorithm (GA) [14, 28], a machine learning method inspired by natural evolution. With sufficient computational resources, this algorithm can reconstruct any network topology, and in our case it has the ability to reconstruct the topology of networks up to 12 nodes with high fidelity. Furthermore, by altering the probing procedure to consider injecting the excitation into different nodes, we can simplify the reconstruction procedure and enable a significant reduction in the required computational power. This facilitates reconstruction of topologies for larger networks when constrained by the available computational resources.
@@
The remainder of this manuscript is structured as follows. In Sec. II we introduce the physical model under study. In Sec. III we discuss the GA used to infer the network’s topology. Finally, in Sec. IV we perform an analysis to determine the optimal algorithm configuration, and report on its performance. Sec. V is devoted to a summary of the main results that we have achieved an provides a further look into interesting directions for future investigation.
@@ II The Model
which describes the transfer of energy through a quantum network of coupled qubits, with $\sigma_{i}^{\pm}$ the ladder operators of qubit $i$. The first term in Eq. (1) represents the on site energy of node $i$, where we assume each qubit has the same on-site energy $\omega$. The second term represents how the excitation hops from one node to another with hopping strength $g_{ij}$, which we assume for simplicity to only take values 0 and 1. We consider the case where only one excitation is traversing the network, thus restricting the dynamics to the single excitation subspace, described by the single excitation Hamiltonian $H^{(1)}$ with elements
@@
where $\{|i\rangle:0\leq i<n\}$ is the basis for the single excitation subspace with $|i\rangle$ representing the physical configuration where the excitation localised at node $i$. Notably, $H^{(1)}$ significantly reduces the complexity of computation, as the dimension of the Hilbert space now scales linearly with the number of nodes, rather than exponentially. As the on-site operators commute with the total Hamiltonian, they do not contribute to the dynamical evolution of the system. We thus define the adjacency matrix $A$ as
@@
with $\gamma$ the decay rate. This mechanism mirrors the function of the reaction centre in the FMO complex, where absorbed light is irreversibly converted into chemical energy. In the model, once the excitation reaches node $s$, it exponentially decays into the sink at a rate controlled by $\gamma$, and cannot re-enter the network once absorbed.
The initial state of the network is chosen as a single excitation injected at the $e^{\text{th}}$ node, and we denote as $\rho(t)$ the state of the system at time $t$, evolving under the Hamiltonian in Eq. (5). The non-Hermiticity of $H_{sink}$ results in dynamics that are no longer excitation-preserving, so that $\rho(t)$ is no longer of trace one. The population loss of the network is accounted for by the increase in the sink population, so that the sum of the two is always one. We have therefore that the population of the sink at time $t$ is given by
@@
$P_{s}(t)\approx\gamma A_{es}t^{3}-\left[\gamma A_{e,s}+\sum_{i}\langle i|A^{2}|e\rangle^{2}\right]\frac{t^{4}}{4}.$ (9)
@@ III The Genetic Algorithm
GAs allow for the optimisation of problems with complex high-dimensional search spaces. A GA utilises iterative update methods inspired by the evolutionary principles of selection, crossover and mutation on candidate solutions. In particular, GAs excel at optimisation problems with discretised search spaces, where traditional optimisation techniques such as stochastic gradient descent are not applicable [28].
@@
In our work, the target and population networks are chosen as $n$-node Erdös-Rényi graphs [29], where each edge connection is included with some probability, independent from other edges. Furthermore, we consider only connected graphs. The sink population data for every network is collected according to the following procedure:
@@
As discussed in Sec. II, it is possible to infer, using the short-time sink population, whether there is a direct connection between nodes $e$ and $s$, regardless of the location of the latter. We thus set known adjacency matrix elements to match that of the target for all members of the initial population and offspring generated by the GA. This reduces the number of optimisation parameters from ${n(n-1)}/{2}$ to ${(n-1)(n-2)}/{2}$.
@@
Figure 2: Flow chart of GA. 1. Initial population and target individual generated. 2. Evaluation of fitness function and ordering of individuals from best to worst. Either the algorithm finds the target network and stops, or weaker individuals are eliminated and stronger individuals selected for genetic evolution. 3. Mating procedure, two selected individual contribute chromosomes to form offspring. 4. Slight genetic mutation are induced in weaker parent and all offspring, adding or removing chromosomes from the population, potentially resulting in new characteristics. The new generation has their fitness evaluated and the algorithm loops until a target solution is found or the algorithm reaches the pre-set iteration limit.
The parents then undergo the mating procedure, in order to restore the original population size. Each parent mates with four randomly selected parents, so that the population is restored to its original size. The mating of parents $a$ and $b$ proceeds as follows
@@ IV.1 Optimal Model Parameters and Performance
This metric describes the total number of topological differences between the networks, ranging from 0 to $n(n-1)/2$, and thus serves an effective measure of distance between them. Using this metric to compute the FDC, it is possible to estimate the reconstructability of a given network topology by means of a GA. As a result, to maximize the performance of the algorithm, it is necessary to choose the times at which measurements on the sink are performed so as to maximize the average FDC of the problem over the space of possible target networks.
@@
Figure 3: Dynamics of FDC over time for fully random Erdös-Rényi networks. Main panel: We show the time at which the maximum value of the FDC is achieved against the number of nodes in a network. The (blue) crosses represent simulated data, while the red line represents an inverse exponential fit. Inset: FDC as a function of time for varying network dimension.
As it is possible to infer whether the excitation and sink nodes are connected, the FDC dynamics in Fig. 3 do not exactly reflect that of the GA setup. When running the algorithm, these connections will always match between the target and the population. As a consequence, it is necessary to consider an alternative setup for the analysis of the FDC. For each target network, a set of trial solution networks were generated with matching adjacency matrix elements where the elements can be inferred from the short-time sink population. Based on this setup, the main panel of Fig. 4 shows the time of at which FDC achieves its maximum as a function of the number of nodes, with the inset showing the FDC as a function of time for different network sizes. In contrast to Fig. 3, the FDC has both smaller peak values and reaches maximal correlation at a later time than in the previous case. This is due to the non-linear relationship between the sink population and the adjacency matrix elements that do not directly connect the sink and the excitation nodes, which become relevant at longer time scales. Once again, the time of peak correlation follows an inverse exponential relationship with system size. Based on this analysis, the sink population will be measured at times centred on the time of maximum FDC.
@@
Figure 5: Convergence of the GA for different network sizes. Main panel: Convergence rate as a function of the population size for varying network sizes. The black arrow shows the direction of growth of $n$. Inset: Number of instances of convergence of the algorithm to the exact target topology for $n=\{6,7,8,9,10\}$ node networks and population sizes $\{100,500,1000,2000,10000,50000\}$ respectively.
Using the previously discussed parameter configuration, the algorithm performance was assessed by estimating the topology of 5000 target networks for system sizes between 6 and 10 nodes. The rate of successful convergence to the target topology is shown as a function of system size in the inset figure of Fig. 5 with population sizes $\{100,500,1000,2000,10000,50000\}$ respectively. The model is capable of correctly reconstructing all target networks up to a size of 8 nodes and achieved a performance of over 97% for $n=9$ and 10.
@@ IV.2 Reconstruction of Larger Networks Using Multiple Injection Locations
Through this probing procedure, we can investigate the necessary number of measurements required to achieve optimal performance for a given population size. In Fig. 6 the performance of the algorithm as a function of the number of probing configurations used is shown for various values of $n$, up to 14 , with a population of 50 networks. The convergence rate was estimated using 1000 sample targets represented by the crosses in the figure. A generalised logistic curve represents an accurate fitting of the data: for all network sizes the convergence rate saturates to 100% given a sufficiently large number of initial measurement configurations. This is to be expected, because if all possible probing configurations are measured, then the structure of the adjacency matrix is explicitly identified by the measurement data.
Figure 6: Convergence rate of the GA studied against the number of measurement configurations as the number $n$ of nodes of a network grows. Simulated data are shown as coloured crosses, while dashed lines represent a generalised logistic fitting of the data. The black arrow shows the direction of growth of $n$.
@@
Figure 7: Average population size required for convergence of the algorithm at a given number of measurement configurations for varying system sizes. Simulated data is marked with crosses and dashed lines represent a first order rational function fitting of the data. The black arrow shows the direction of growth of $n$.
The ability of the algorithm to reconstruct a given network is clearly highly dependent on the size of population and the number of measurement configurations utilized during the measurement procedure. This is displayed in Fig. 7, which shows the population size needed to guarantee the reconstruction of a given network as a function of the number of measurement configurations being used. By increasing the latter, the necessary population size for network reconstruction decreases, and so becomes more feasible for simulation on computers with limited resources. This highlights the necessary trade-off between effort invested in collection of sink data, depending on the number of measurement configurations, and the computational resources that are available for the problem.
@@
Figure 8: Average number of unknown network connections that can be successfully reconstructed as a function of population size in the algorithm for various sizes of target network. Simulated data is marked with crosses and dashed lines represent a first order rational function fitting of the data. The black arrow shows the direction of growth of $n$
@@ V Conclusions
We have investigated the ability of GAs to reconstruct the topology of a quantum network using indirect probe-based measurements. While the proposed methodology is entirely general, we have benchmarked its performance by demonstrating that it can effectively reconstruct networks of up to 12 nodes, with performance on larger system sizes limited by the computational requirements of simulating a sufficiently large population.
@@ Acknowledgements.
C.J.C. and M.M. are grateful to the Quantum Theory group at University of Palermo (https://quantum.unipa.it/) for hospitality during various stages of development of this project. We acknowledge support from the European Union’s Horizon Europe EIC-Pathfinder project QuCoM (101046973), the Department for the Economy of Northern Ireland under the US-Ireland R&D Partnership Programme, the “Italian National Quantum Science and Technology Institute (NQSTI)" (PE0000023) - SPOKE 2 through project ASpEQCt, the “National Centre for HPC, Big Data and Quantum Computing (HPC)” (CN00000013) – SPOKE 10 through project HyQELM, and the EU Horizon Europe EIC Pathfinder project QuCoM (GA no. 10032223), the Italian Ministry of University and Research under PNRR - M4C2-I1.3 Project PE-00000019 "HEAL ITALIA" (CUP B73C22001250006). SD acknowledges support from Istituto Nazinale di Fisica Nucleare (INFN).

### v3

@@ I Introduction
Quantum networks provide a powerful platform for distributing, processing, and storing quantum information across multiple interconnected systems, playing a significant role in the advancement of emerging quantum technologies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. Reconstructing the topology of unknown networks is therefore a fundamental task for quantum technological applications. Continuous-time quantum walks (CTQWs) [14, 15] are an effective tool for modelling the topology-driven dynamics in complex quantum networks. They have been successfully used to investigate the dynamics of transport phenomena in numerous systems [16, 17]. A key example is the demonstration of efficient energy transfer in the Fenna-Matthews-Olson complex [18], one of the most extensively studied light-harvesting protein-pigment complexes [19, 20, 21]. The complete characterisation of a network’s topology can be achieved using methods such as quantum process tomography [22], which requires however direct measurement of the state of the entire network, with the number of required measurements scaling exponentially with the system’s dimension. As a result, such an approach quickly becomes infeasible for even moderate network sizes [23].
@@
This work tackles a fundamental challenge of inferring the topology of an unknown quantum network [24] by leveraging indirect system measurements using a probe. Approaches based on direct measurements, such as  [14] provide alternative methods for topology reconstruction for up to 10 node networks with noisy direct measurements. The novelty of our approach is that it provides a non-invasive framework for interacting with complex quantum systems without the need for direct access to the network itself. Such probing procedure consists of injecting an excitation into a network node and attaching an external sink to another one, which we model as an irreversible decay channel. The excitation crosses the network to decay into the sink. By measuring the sink, we collect data driven by the underlying network topology. We thus seek to exploit this relationship to infer the topology of the network.
@@
We frame the task as an optimization problem with the goal of constructing a network with measurement data as similar to the target network as possible. As the parameters of this problem are the connections between network nodes, the parameter space is discretised and so lends itself to an approach such as a genetic algorithm (GA) [14, 25], a machine learning method inspired by natural evolution. With sufficient computational resources, this algorithm can reconstruct any network topology, and in our case it has the ability to reconstruct the topology of networks up to 10 nodes with high fidelity. Furthermore, by altering the probing procedure to consider injecting the excitation into different nodes, we can simplify the reconstruction procedure and enable a significant reduction in the required computational power. This facilitates reconstruction of topologies for larger networks when constrained by the available computational resources.
@@
The remainder of this manuscript is structured as follows. In Sec. II we introduce the physical model under study. In Sec. III we discuss the GA used to infer the network’s topology. Finally, in Sec. IV we perform an analysis to determine the optimal algorithm configuration, and report on its performance. Sec. V is devoted to a summary of the main results that we have achieved and provides a further look into interesting directions for future investigation.
@@ II The Model
which describes the transfer of energy through a quantum network of coupled qubits, with $\sigma_{i}^{\pm}$ the ladder operators of qubit $i$. The first term in Eq. (1) represents the on-site energy of node $i$, where we assume each qubit has the same on-site energy $\omega$. The second term represents how the excitation hops from one node to another with hopping strength $g_{ij}$, which we assume for simplicity to only take values 0 and 1. We consider the case where only one excitation is traversing the network, thus restricting the dynamics to the single excitation subspace, described by the single excitation Hamiltonian $H^{(1)}$ with elements
@@
where $\{|i\rangle:0\leq i<n\}$ is the basis for the single excitation subspace with $|i\rangle$ representing the physical configuration where the excitation is localised at node $i$. Notably, $H^{(1)}$ significantly reduces the complexity of computation, as the dimension of the Hilbert space now scales linearly with the number of nodes, rather than exponentially. As the on-site operators commute with the total Hamiltonian, they do not contribute to the dynamical evolution of the system. We thus define the adjacency matrix $A$ as
@@
with $\gamma$ the decay rate. This mechanism mirrors the function of the reaction centre in the FMO complex, where absorbed light is irreversibly converted into chemical energy. In the model, once the excitation reaches node $s$, it exponentially decays into the sink at a rate controlled by $\gamma$, and cannot re-enter the network once absorbed. A schematic of the network and external sink probe is shown in Fig. 1
Figure 1: Pictorial representation of a quantum network of interconnected nodes, with the excitation (purple star) being initially localised in one node, and an externally attached sink which is used to perform indirect measurements (green eye)
The initial state of the network is chosen as a single excitation injected at the $e^{\text{th}}$ node, and we denote $\rho(t)$ as the state of the system at time $t$, evolving under the Hamiltonian in Eq. (5). The non-Hermiticity of $H_{sink}$ results in dynamics that are no longer excitation-preserving, so that $\rho(t)$ is no longer of trace one. The population loss of the network is accounted for by the increase in the sink population, so that the sum of the two is always one. We have therefore that the population of the sink at time $t$ is given by
@@
$P_{s}(t)\approx\gamma A_{es}t^{3}-\left[\gamma A_{es}+\sum_{i}\langle i|A^{2}|e\rangle^{2}\right]\frac{t^{4}}{4}.$ (9)
@@ III The Genetic Algorithm
GAs allow for the optimisation of problems with complex high-dimensional search spaces. A GA utilises iterative update methods inspired by the evolutionary principles of selection, crossover, and mutation on candidate solutions. In particular, GAs excel at optimisation problems with discretised search spaces, where traditional optimisation techniques such as stochastic gradient descent are not applicable [25].
@@
In our work, the target and population networks are chosen as $n$-node Erdös-Rényi graphs [26], where each edge connection is included with random probability $p\in[0,1]$, resulting in networks of all different edge densities. Furthermore, we consider only connected graphs. The sink population data for every network is collected according to the following procedure:
@@
As discussed in Sec. II, it is possible to infer, using the short-time sink population, whether there is a direct connection between nodes $e$ and $s$, regardless of the location of the latter. We thus set known adjacency matrix elements to match that of the target for all members of the initial population and offspring generated by the GA. This reduces the number of optimisation parameters from n(n1)/2{n(n-1)}/{2} to (n1)(n2)/2{(n-1)(n-2)}/{2}.
@@
Figure 3: Flow chart of GA. 1. Initial population and target individual generated. 2. Evaluation of fitness function and ordering of individuals from best to worst. Either the algorithm finds the target network and stops, or weaker individuals are eliminated and stronger individuals selected for genetic evolution. 3. Mating procedure, two selected individuals contribute chromosomes to form offspring. 4. Slight genetic mutations are induced in weaker parent and all offspring, adding or removing chromosomes from the population, potentially resulting in new characteristics. The new generation has their fitness evaluated and the algorithm loops until a target solution is found or the algorithm reaches the pre-set iteration limit.
The parents then undergo the mating procedure, in order to restore the original population size. Each parent mates with four randomly selected parents. The mating of parents $a$ and $b$ proceeds as follows
@@ IV.1 Optimal Model Parameters and Performance
This metric describes the total number of topological differences between the networks, ranging from 0 to n(n1)/2n(n-1)/2, and thus serves an effective measure of distance between them. Using this metric to compute the FDC, it is possible to estimate the reconstructability of a given network topology by means of a GA. As a result, to maximize the performance of the algorithm, it is necessary to choose the times at which measurements on the sink are performed so as to maximize the average FDC of the problem over the space of possible target networks.
@@
Figure 4: Dynamics of FDC over time for fully random Erdös-Rényi networks. Main panel: We show the time at which the maximum value of the FDC is achieved against the number of nodes in a network. The blue crosses represent simulated data, while the red line represents an inverse exponential fit. Inset: FDC as a function of time for varying network dimension.
As it is possible to infer whether the excitation and sink nodes are connected, the FDC dynamics in Fig. 4 do not exactly reflect that of the GA setup. When running the algorithm, these connections will always match between the target and the population. As a consequence, it is necessary to consider an alternative setup for the analysis of the FDC. For each target network, a set of trial solution networks were generated with matching adjacency matrix elements where the elements can be inferred from the short-time sink population. Based on this setup, the main panel of Fig. 5 shows the time at which FDC achieves its maximum as a function of the number of nodes, with the inset showing the FDC as a function of time for different network sizes. In contrast to Fig. 4, the FDC has both smaller peak values and reaches maximal correlation at a later time than in the previous case. This is due to the non-linear relationship between the sink population and the adjacency matrix elements that do not directly connect the sink and the excitation nodes, which become relevant at longer time scales. Once again, the time of peak correlation follows an inverse exponential relationship with system size. Based on this analysis, the sink population will be measured at times centred on the time of maximum FDC.
@@
Figure 6: Convergence of the GA for different network sizes. Main panel: Convergence rate as a function of the population size for varying network sizes. The black arrow shows the direction of growth of $n$. Inset: Number of instances of convergence of the algorithm to the exact target topology for $n=\{6,7,8,9,10\}$ node networks and population sizes $\{100,500,1000,2000,10000\}$ respectively.
Using the previously discussed parameter configuration, the algorithm performance was assessed by estimating the topology of 5000 target networks for system sizes between 6 and 10 nodes. The rate of successful convergence to the target topology is shown as a function of system size in the inset of Fig. 6 with population sizes $\{100,500,1000,2000,10000\}$ respectively. The model is capable of correctly reconstructing all target networks up to a size of 8 nodes and achieved a performance of over 97% for $n=9$ and 10. Networks that the algorithm failed to reconstruct were predominantly of a medium connectivity, centered around the area of the search space with a large number of permutations for a given number of connections. As a result, the model is likely to perform better for networks that are either sparsely or densely connected.
The model was ran on a laptop utilising a 10-core CPU with a 3.5GHz clock-speed and 16GB RAM. For networks of 10 nodes, the algorithm had an average runtime per iteration of approximately 2.3s, with an average of 25.5 iterations leading to a total average runtime of roughly 58s. The algorithm utilised roughly 2GB of working memory. The runtime and memory requirements for the algorithm scales close to linearly with the number of network nodes fixing the population size. The limiting factor therefore, is the exponentially increasing population size necessary to reconstruct a target network as the number of nodes is increased. Increasing memory requirements can be handled by batch processing portions of the population sequentially, and so the limiting factor rapidly becomes the runtime of the algorithm for larger system sizes.
@@ IV.2 Reconstruction of Larger Networks Using Multiple Injection Locations
Through this probing procedure, we can investigate the necessary number of measurements required to achieve optimal performance for a given population size. In Fig. 7 the performance of the algorithm as a function of the number of probing configurations used is shown for various values of $n$, up to 14 , with a population of 50 networks. The convergence rate was estimated using 1000 sample targets represented by the crosses in the figure. A generalised logistic curve represents an accurate fitting of the data: for all network sizes the convergence rate saturates to 100% given a sufficiently large number of initial measurement configurations. This is to be expected, because if all possible probing configurations are measured, then the structure of the adjacency matrix is explicitly identified by the measurement data. As a result, the problem of reconstructing large networks is reduced to collecting enough measurement data to reduce the size of the problem to that which can be achieved given computational resources.
Figure 7: Convergence rate of the GA studied against the number of measurement configurations as the number $n$ of nodes of a network grows. Simulated data is represented with coloured markers, while dashed lines represent a generalised logistic fitting of the data. The black arrow shows the direction of growth of $n$.
@@
Figure 8: Average population size required for convergence of the algorithm at a given number of measurement configurations for varying system sizes. Simulated data is represented with coloured markers and dashed lines represent a first order rational function fitting of the data. The black arrow shows the direction of growth of $n$.
The ability of the algorithm to reconstruct a given network is clearly highly dependent on the size of population and the number of measurement configurations utilized during the measurement procedure. This is displayed in Fig. 8, which shows the population size needed to guarantee the reconstruction of a given network as a function of the number of measurement configurations being used. By increasing the latter, the necessary population size for network reconstruction decreases, thus providing a method to mitigate the exponential increase in computational resources required as system size increases (Fig. 6), and making this approach more feasible for simulation on computers with limited resources. This highlights the necessary trade-off between effort invested in collection of sink data, depending on the number of measurement configurations, and the computational resources that are available for the problem. As a result, this method facilitates the reconstruction of any size of network given a number of measurements and a given population size. However, the necessary computational resources and experimental requirements will likely become very demanding for very large network sizes.
@@
Figure 9: Average number of unknown network connections that can be successfully reconstructed as a function of population size in the algorithm for various sizes of target network. Simulated data represented by coloured markers and dashed lines represent a first order rational function fitting of the data. The black arrow shows the direction of growth of $n$
@@ V Conclusions
We have investigated the ability of a GA to reconstruct the topology of a quantum network using indirect probe-based measurements. While the proposed methodology is entirely general, we have benchmarked its performance by demonstrating that it can effectively reconstruct networks of up to 10 nodes, with performance on larger system sizes limited by the computational requirements of simulating a sufficiently large population.
@@ Acknowledgements.
C.J.C. and M.M. are grateful to the Quantum Theory group at University of Palermo (https://quantum.unipa.it/) for hospitality during various stages of development of this project. We acknowledge support from the European Union’s Horizon Europe EIC-Pathfinder project QuCoM (101046973), the Department for the Economy of Northern Ireland under the US-Ireland R&D Partnership Programme, the “Italian National Quantum Science and Technology Institute (NQSTI)" (PE0000023) - SPOKE 2 through project ASpEQCt.
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.